# Not All Samples Should Be Utilized Equally: Towards Understanding and Improving Dataset Distillation

### v1

@@ 1 Introduction
In an era of data-centric AI, scaling laws [19] have shifted the focus to data quality. Under this scenario, dataset distillation (DD) [46, 39] has emerged as a solution for creating high-quality data summaries. Unlike data pruning methods [13, 7, 45, 1] that directly select data points from original datasets, DD methods are designed to generate novel data points through learning. The utility of DD methods has been witnessed in fields such as privacy protection [6, 29, 4, 11], continual learning [38, 15, 50, 32], neural architecture search [2, 34, 43], and federated learning [14, 28, 18].
@@
To address Question 1, we first empirically examine the differences between matching-based distillation methods. It is widely acknowledged that sample difficulty (Definition 1) is a crucial metric in data-centric AI that significantly affects model performance, as seen in dataset pruning [44, 31, 30, 42], active learning [40, 5, 49], and large language model prediction [9, 27, 26]. To track the differences between current distillation methods, we follow [37] and analyze sample difficulty using the GraDN metric (Definition 2). Surprisingly, we discover that the GraDN score is increased in GM-based methods (Figure 1(a)), while TM-based methods may reduce this metric (Figure 1(c)). These distinct trends indicate that the difficulty of samples utilized in GM-based methods is elevated (Figure 2(a)), whereas in TM-based methods, it is reduced (Figure 2(b)) during the distillation process.
Motivated by these empirical observations, we develop a theoretical explanation for current DD methods from the perspective of sample difficulty. Specifically, we draw upon the neural scaling law in the data pruning theory [42] to connect sample difficulty with distillation performance. As shown in Figure 4(c), our theory indicates that in matching-based DD methods, when the synthetic dataset is small—specifically, when the images-per-class (IPC) is low—the optimal strategy is to primarily focus on easier samples rather than harder ones to enhance data quality. Based on our theory, we further explain why TM-based methods usually outperform GM-based methods in real scenarios.
Beyond developing a theoretical framework, we take further steps to seek possible solutions for enhancing current approaches. This introduces another important research question:
@@
To address Question 2, inspired by our empirical observations and theoretical analysis, we propose our novel Sample Difficulty Correction (SDC) method to enhance the synthetic dataset quality learned by current matching-based distillation methods. We achieve this by imposing the distillation method to focus more on learning from easy samples rather than hard samples. Specifically, we implicitly add a gradient norm regularizer to improve the synthetic dataset quality.
@@
  • We empirically investigate sample difficulty from the gradient norm perspective across various distillation methods, establishing a link between sample difficulty and synthetic dataset quality. We propose that GM-based methods tend to learn difficult samples during optimization, whereas TM-based methods do not exhibit a dominant preference for difficulty. This difference may explain the poorer performance of GM-based methods than TM-based methods in real-world applications.
  • We theoretically elucidate the mechanism of matching-based DD methods from the perspective of sample difficulty with statistical mechanics. Adapting the neural scaling law theory from data pruning [42] to distillation settings, we provide insights into how matching strategies evolve with the size of the synthetic dataset. Consequently, we propose that focusing on matching easy samples is a better strategy when the synthetic dataset is small.
  • We introduce Sample Difficulty Correction (SDC) to improve the quality of synthetic datasets in current matching-based DD methods. SDC can be incorporated into existing methods with minimal code modifications. Our method demonstrates superior generalization performance across 7 distillation methods (DC [52], DSA [51], DSAC [25], MTT [3], FTD [12], TESLA [8], and DATM [16]) and 6 datasets (MNIST [10], FashionMNIST [48], SVHN [36], CIFAR-10 [21], CIFAR-100 [21], and Tiny ImageNet [23]).
@@ 2 Preliminaries and Related Work
Gradient Matching (GM) based methods are pivotal in achieving distillation by ensuring the alignment of training gradients between surrogate models trained on both the original dataset $\mathcal{D}_{\mathsf{real}}$ and the synthesized dataset $\mathcal{D}_{\mathsf{syn}}$. This method is first introduced by DC [52]. Let $\theta_{t}$ represent the network parameters sampled from distribution $P_{\theta}$ at step $t$, and $C$ symbolizes the categories within $\mathcal{D}_{\mathsf{real}}$. The cross-entropy loss, denoted by $\mathcal{L}$, is employed to assess the matching loss by comparing the gradient alignment over a time horizon of $T$ steps. The formal optimization objective of DC is:
$\underset{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}{\operatorname{\arg\min}}\underset{\theta_{0}\sim P_{\theta},c\sim C}{\mathbb{E}}\left[\sum_{t=0}^{T}\mathbf{D}\left(\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}^{c}}\left(\theta_{t}\right),\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}^{c}}\left(\theta_{t}\right)\right)\right],$ (1)
@@
$\theta_{t+1}\leftarrow\theta_{t}-\eta\cdot\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}\left(\theta_{t}\right).$ (2)
@@
$\underset{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}{\operatorname{\arg\min}}\underset{\theta_{0}\sim P_{\theta}}{\mathbb{E}}\left[\sum_{t=0}^{T}\mathbf{D}\left(\mathbb{E}_{c\in C}\left[\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}^{c}}(\theta_{t})\right],\mathbb{E}_{c\in C}\left[\nabla_{\theta}\mathcal{L}_{\mathcal{D}^{c}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}(\theta_{t})\right]\right)\right].$ (3)
@@
$\underset{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}{\arg\min}\underset{\theta_{0}\sim P_{\theta}}{\mathbb{E}}\left[\sum_{t=0}^{T-M}\frac{\mathbf{D}\left(\theta_{t+M}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}},\theta_{t+N}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}\right)}{\mathbf{D}\left(\theta_{t+M}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}},\theta_{t}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}}\right)}\right],$ (4)
@@
${\theta}_{t+i+1}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}\leftarrow{\theta}_{t+i}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}-\eta\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}({\theta}_{t+i}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}),\text{ where }{\theta}_{t}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}:=\theta_{t}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}}.$ (5)
Along this line, TESLA [8] utilizes linear algebraic manipulations and soft labels to increase compression efficiency, FTD [12] aims to seek a flat trajectory to avoid accumulated trajectory error, and DATM [16] considers matching only necessary parts of trajectory with difficulty alignment.
@@ 3.1 A Closer Look at Sample Difficulty
Empirical verification of the relationship between sample difficulty and gradient norm. We conducted experiments to verify the reliability of the GraDN score in classifying the CIFAR-10 and CIFAR-100 datasets by training a set of models. As depicted in Figure 3, the GraDN score shows a clear positive correlation with the sample difficulty. For easier samples, GraDN scores are generally lower, exerting minimal impact on the network’s gradient flow. Conversely, for harder samples, higher GraDN scores indicate a significant impact on the optimization directions of the models. We show detailed results of the relationships between these metrics in Appendix A.
@@ 3.2 An Analytical Theory for Explaining Matching-based Dataset Distillation
We analyze the test error $\varepsilon$ of this final perceptron $\theta^{\mathcal{D}_{\mathsf{syn}}}$ as a function of $\alpha_{\mathsf{syn}}$, $f$, and the angle $\gamma$ between the probe student $\theta^{\mathsf{probe}}$ and the expert $\theta^{\mathcal{D}_{\mathsf{real}}}$. Similar to [42], we define $\theta^{\mathsf{probe}}$ as a random Gaussian vector conditioned to have angle $\gamma$ with the expert. Under this scenario, we can derive an asymptotically exact analytic expression for $\varepsilon(\alpha_{\mathsf{syn}},f,\gamma)$ (see Appendix D for details):
$\varepsilon=\frac{\cos^{-1}(R)}{\pi},\quad\text{where }R=\frac{\theta^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}^{\top}}\theta^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}}{\|\theta^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\mathsf{real}}}}\|_{2}\|\theta^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\mathsf{syn}}}}\|_{2}}$ (8)
Likewise in [42], we can solve $R$ with saddle point equations in Appendix D.1, enabling direct predictions of the test error $\varepsilon$ according to Eq (8).
Verification of the neural scaling law in dataset distillation. We first evaluated the correctness of our theory in the perfect expert-student setting ($\gamma=0$). As shown in Figure 4(a), we observed an excellent match between our analytic theory (solid curves) and numerical simulations (dots) of perceptron learning at parameters $d=200$ in dataset distillation. We also verify our theory in imperfect probe settings when $\gamma\neq 0$, as shown in Appendix D.2.

### v2

@@ 1 Introduction
In an era of data-centric AI, scaling laws [17] have shifted the focus to data quality. Under this scenario, dataset distillation (DD) [45, 36, 44, 43] has emerged as a solution for creating high-quality data summaries. Unlike data pruning methods [13, 7, 41, 1, 42, 49] that directly select data points from original datasets, DD methods are designed to generate novel data points through learning. The utility of DD methods has been witnessed in fields such as privacy protection [6, 26, 4, 11], continual learning [35, 14, 50, 29], and neural architecture search [2, 31, 39]
@@
To address Question 1, we first empirically examine the differences between matching-based distillation methods. It is widely acknowledged that sample difficulty (Definition 1) is a crucial metric in data-centric AI that significantly affects model performance, as seen in dataset pruning [40, 28, 27, 38], and large language model prediction [9, 25, 24]. To track the differences between current distillation methods, we follow [34] and analyze sample difficulty using the GraDN metric (Definition 2). Surprisingly, we discover that the GraDN score is increased in GM-based methods (Figure 1(a)), while TM-based methods may reduce this metric (Figure 1(c)). These distinct trends indicate that the difficulty of samples utilized in GM-based methods is elevated (Figure 2(a)), whereas in TM-based methods, it is reduced (Figure 2(b)) during the distillation process.
Motivated by these observations, we develop a theoretical explanation for current DD methods from the perspective of sample difficulty. Specifically, we draw upon the neural scaling law in the data pruning theory [38] to connect sample difficulty with performance. As shown in Figure 4(c), our theory indicates that in matching-based DD methods, when the synthetic dataset is small—specifically, when the images-per-class (IPC) is low—the optimal strategy is to primarily focus on easier samples rather than harder ones to enhance performance. Based on our theory, we further explain why TM-based methods usually outperform GM-based methods in real scenarios.
Beyond developing a theoretical framework, we take steps to explore solutions for improving current approaches. This raises another key research question:
@@
To address Question 2, based on our empirical observations and theoretical analysis, we propose the novel Sample Difficulty Correction (SDC) method to improve the synthetic dataset quality in current matching-based distillation methods. We do this by guiding the distillation method to focus more on easy samples than hard samples, adding an implicit gradient norm regularizer to enhance quality.
@@
  • We empirically investigate sample difficulty from the perspective of gradient norm in distillation methods, linking it to synthetic dataset quality. We propose that GM-based methods focus on difficult samples during optimization, while TM-based methods show no dominant preference for difficulty. This may explain the poorer performance of GM-based methods than TM-based methods.
  • We theoretically elucidate the mechanism of matching-based DD methods from the perspective of sample difficulty. Adapting the neural scaling law theory from data pruning [38] to distillation settings, we provide insights into how matching strategies evolve with the size of the synthetic dataset. Consequently, we propose that focusing on matching easy samples is a better strategy when the synthetic dataset is small.
  • We introduce Sample Difficulty Correction (SDC) to improve the quality of synthetic datasets in current matching-based DD methods. Our method demonstrates superior generalization performance across 7 distillation methods (DC [52], DSA [51], DSAC [23], MTT [3], FTD [12], TESLA [8], DATM [15]) and 6 datasets (MNIST [10], FashionMNIST [47], SVHN [33], CIFAR-10/100 [19], and Tiny-ImageNet [21]).
@@ 2 Preliminaries and Related Work
Gradient Matching (GM) based methods are pivotal in achieving distillation by ensuring the alignment of training gradients between surrogate models trained on both the original dataset $\mathcal{D}_{\mathsf{real}}$ and the synthesized dataset $\mathcal{D}_{\mathsf{syn}}$. This method is first introduced by DC [52]. Let $\theta_{t}$ represent the network parameters sampled from distribution $P_{\theta}$ at step $t$, and $C$ symbolizes the categories within $\mathcal{D}_{\mathsf{real}}$. The cross-entropy loss $\mathcal{L}$, is employed to assess the matching loss by comparing the gradient alignment over a time horizon of $T$ steps. The formal optimization objective of DC is:
$\underset{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}{\operatorname{\arg\min}}\underset{\theta_{0}\sim P_{\theta},c\sim C}{\mathbb{E}}\left[\sum_{t=0}^{T}\mathbf{D}\left(\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}^{c}}\left(\theta_{t}\right),\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}^{c}}\left(\theta_{t}\right)\right)\right],$ (1)
@@
$\theta_{t+1}\leftarrow\theta_{t}-\eta\cdot\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}\left(\theta_{t}\right).$ (2)
@@
$\begin{split}\underset{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}{\operatorname{\arg\min}}\underset{\theta_{0}\sim P_{\theta}}{\mathbb{E}}\left[\sum_{t=0}^{T}\mathbf{D}\left(\mathbb{E}_{c\in C}\left[\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}^{c}}(\theta_{t})\right],\right.\right.\\ \left.\left.\mathbb{E}_{c\in C}\left[\nabla_{\theta}\mathcal{L}_{\mathcal{D}^{c}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}(\theta_{t})\right]\right)\right].\end{split}$ (3)
@@
$\underset{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}{\arg\min}\underset{\theta_{0}\sim P_{\theta}}{\mathbb{E}}\left[\sum_{t=0}^{T-M}\frac{\mathbf{D}\left(\theta_{t+M}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}},\theta_{t+N}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}\right)}{\mathbf{D}\left(\theta_{t+M}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}},\theta_{t}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}}\right)}\right],$ (4)
@@
${\theta}_{t+i+1}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}\leftarrow{\theta}_{t+i}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}-\eta\nabla_{\theta}\mathcal{L}_{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}({\theta}_{t+i}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}),\text{ where }{\theta}_{t}^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}:=\theta_{t}^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}}.$ (5)
Similarly, TESLA [8] utilizes linear algebraic manipulations and soft labels to increase compression efficiency, FTD [12] aims to seek a flat trajectory to avoid accumulated trajectory error, and DATM [15] considers matching only necessary parts of trajectory with difficulty alignment.
@@ 3.1 A Closer Look at Sample Difficulty
Empirical verification of the relationship between sample difficulty and gradient norm. We conducted experiments to verify the reliability of the GraDN score in classifying the CIFAR-10 and CIFAR-100 datasets by training a set of models. As depicted in Figure 3, the GraDN score shows a clear positive correlation with the sample difficulty. For easier samples, GraDN scores are generally lower, exerting minimal impact on the network’s gradient flow. Conversely, for harder samples, higher GraDN scores indicate a significant impact on the optimization directions of the models. We show detailed results of the relationships between these metrics in Appendix 1.
@@ 3.2 An Analytical Theory for Explaining Matching-based Dataset Distillation
We analyze the test error $\varepsilon$ of this final perceptron $\theta^{\mathcal{D}_{\mathsf{syn}}}$ as a function of $\alpha_{\mathsf{syn}}$, $f$, and the angle $\gamma$ between the probe student $\theta^{\mathsf{probe}}$ and the expert $\theta^{\mathcal{D}_{\mathsf{real}}}$. Similar to [38], we define $\theta^{\mathsf{probe}}$ as a random Gaussian vector conditioned to have angle $\gamma$ with the expert. Under this scenario, we can derive an asymptotically exact analytic expression for $\varepsilon(\alpha_{\mathsf{syn}},f,\gamma)$ (see Appendix 4 for details):
$\varepsilon=\frac{\cos^{-1}(R)}{\pi},\quad\text{where }R=\frac{\theta^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}^{\top}}\theta^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}}{\|\theta^{\mathcal{D}_{{\color[rgb]{1,0,0.5}\mathsf{real}}}}\|_{2}\|\theta^{\mathcal{D}_{{\color[rgb]{0,0.5,1}\mathsf{syn}}}}\|_{2}}$ (8)
Likewise in [38], we can solve $R$ with saddle point equations in Appendix 4.1, enabling direct predictions of the test error $\varepsilon$ according to Eq (8).
Verification of the neural scaling law in dataset distillation. We first evaluated the correctness of our theory in the perfect expert-student setting ($\gamma=0$). As shown in Figure 4(a), we observed an excellent match between our analytic theory (solid curves) and numerical simulations (dots) of perceptron learning at parameters $d=200$ in dataset distillation. We also verify our theory in imperfect probe settings when $\gamma\neq 0$, as shown in Appendix 4.2.